Bakery A uses 300 bags of flour each month. The flour is purchased from a supplier for a price of $100 per bag and an ordering cost of $10 per order. Bakery A’s annual inventory holding cost is 20% of the item value. Bakery A is required to order an integer multiple of 50 bags. What order quantity should Bakery A use to minimize ordering and holding costs? ____ bags (round up to the nearest integer).
Answer #1

DEMAND = 3600
ORDERING COST = 10
HOLDING COST = 20
COST PER UNIT = 100

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 3600 * 10 / 20) = 60

SINCE EOQ = 60, THE NEAREST VALUE THAT IS A MULTIPLE OF 50 WOULD BE 50 AND 100, 50 IS CLOSER TO 60 AND THEREFORE, WE WILL ORDER 50 UNITS TO MINIMIZE THE COST.